40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 21 02:42:17 2014, rana, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00002.PNGTEK00003.PNG
    Reply  Thu Aug 21 19:04:52 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00005.PNG
       Reply  Tue Aug 26 23:16:49 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH Ytfs.pdfpdhComparison.pdfnewY.zip
Message ID: 10430     Entry time: Tue Aug 26 23:16:49 2014     In reply to: 10420
Author: ericq 
Type: Configuration 
Category: Green Locking 
Subject: Gain changes on Green Y PDH 

 Yesterday I measured the spectra and OLTF of the Y-Arm green PDH, after the LO touch-up and 60Hz hunt from last week. I also went to lower frequencies with the SR785, but forgot to take some of the background spectra down there, so I don't have the full breakdown plots yet. Nevertheless, here is the improvement in the PDH error signal:

pdhComparison.pdf

I also measured the OLTF (SR785 injection at the error signal, Auto level ref 5mV at channel 2, 10mV/s source ramping, 50mV max output)

Ytfs.pdf

As you can see, we have tons of phase margin. Flipping the local boost switch had no visible effect on the OLTF; we should change it to something that puts this surplus of phase to good use, and squash the error signal even more. Putting an integrator at 5kHz should still leave about 45 degrees phase margin at 10k. I've started making a LISO model of the PDH board from the DCC drawing, and then I'll inspect the boards individually to make sure I catch the homegrown modifications. 

Data, and code used to generate the plots is attached. 

Attachment 3: newY.zip  412 kB  Uploaded Wed Sep 3 15:05:05 2014
ELOG V3.1.3-